Copper wire has a diameter of 1.02 mm. It carries a current of 1.67 A to a 200 W bulb. The density of free electrons is 8.5x10^28 electrons per cubic meter.

Find a) the current density and b) the drift velocity
Answer #1

a)

current density

J = i/A = 2043742.829 A/m^2

b)

drift velocity

Vd = I/neA

Vd = 0.000150275 m/sec